OpenStudy (anonymous):

(5TH GRADE MATH HELP) To add (30 + 5) + 45, Mark first added 5 and 45. Which property did Mark use to add mentally? A) Identity Property of Addition B) Associative Property of Addition C) Zero Property of Addition D) None

OpenStudy (anonymous):

@ganeshie8 @nincompoop

OpenStudy (anonymous):

@e.mccormick @Zale101 @mathmate

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Associative property

OpenStudy (anonymous):

Okay @texaschic101 do you agree?

OpenStudy (anonymous):

@StudyGurl14 @Jhannybean

OpenStudy (anonymous):

@sleepyjess

OpenStudy (texaschic101):

I agree....associative property

OpenStudy (anonymous):

Ok thank you

OpenStudy (texaschic101):

sure thing :)

OpenStudy (mathmate):

@rishi_k texaschic101 Guys, please explain to the asker why it is the answer you proposed. We're not here to give direct answers and the asker is not learning anything. Please, if you know why your answer is correct, explain it. If you don't, leave the question alone. @Tom_Boy_Rebel Associative property of addition is the property that we can regroup the additions differently and still get the same answer. For example, 1+2+3=(1+2)+3=1+(2+3)=6. We can add 1+2=3 and then 3+3=6, or (2+3)=5 and then add 1 to get 6. See http://www.coolmath.com/prealgebra/06-properties/03-properties-associative-addition-01 for more examples.

OpenStudy (anonymous):

Then @mathmate Associative property states that for any real numbers A,B&C (A+B)+C=A+(B+C) And (A*B)*C=A*(B*C)

OpenStudy (mathmate):

@rishi_K Beautiful! You have also explained associative property on addition and on multiplication... and... thank you!